1/2+(2/3/3/4)-(4/5*5/6)=
Answer by actuary(112)

1/2+(2/3)/(3/4)+3/4*5/6=
1/2+(2/3*4/3)+(3*5)/(4*6)= 1/2+8/9+5/8=(36+84+45)/72 = 165/72= (4*53)/(3*24)=
53/(3*6)=53/18
I hope that this helps.